OpenStudy (anonymous):

Simplify: Cubed root of 16x^5 y^7

OpenStudy (anonymous):

\[\sqrt[b]{x^a} = x^{a/b}\]

OpenStudy (anonymous):

and: \[\sqrt[3]{abc} = \sqrt[3]{a}\sqrt[3]{b}\sqrt[3]{c}\]

OpenStudy (anonymous):

\[\sqrt[3]{16} \sqrt[3]{x^5} \sqrt[3]{y^7}\] Is that what you are saying?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would I simplify \[\sqrt[3]{16}\]